1.      Assume that consumer A consumes two products Q1 and Q2  when the first product price increases, the consumer decided to decrease for the second product? Explain

2.      Given that the profit function for a firm p = 150x + 4x2 – xy – 9y2 + 100y + 2x + y = 50 with a constraint of 2x + y = 50. Determine the amount of x and y which maximizes profit. A. Using the substitution method  B. Using the language method


1
Expert's answer
2021-04-07T10:23:37-0400

1.

the two products are complements. When the price of Q1 increased, the consumer was forced to decrease the consumption of product Q2


2.

A. Substitution method

"y=50-2x"

substituting into the profit function we get

"150x+4x^2-50x-2x^2 -22500+1800=36x^2+5000-200x+2x+50-2x=50"

"38x^2-100x-15700=0"

"x=21.68461851"

"y=6.63076298"


B. Lagrange method

"\frac{\delta f}{\delta x} =\frac{\delta g}{\delta x} \lambda , \frac {\delta f}{\delta y} =\frac {\delta g}{\delta y}\lambda"


"\frac{\delta g}{\delta x}=2\lambda, \frac{\delta g}{\delta y} =\lambda"


"g(x,y) =2x+y=50"

"2(2y)+y=50"

"4y+y=50"

"5y=50"

"\frac{5y}{5}=\frac{50}{5}"

"y=10"


"2x-y=50"

since y=10 we substitute with 10

"2x+10=50"

"2x=50-10"

"2x=40"

"\frac{2x}{2}=\frac{40}{2}"

"x=20"